# Deep Learning based Visually Rich Document Content Understanding: A Survey

### v1

@@ 3. Mono-Task Document Understanding Frameworks
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=3cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2 [Mono-task based models (Content Understanding), for tree=fill=lime!60 [Key Information Extraction, for tree=fill=cyan!40 [Feature-driven Models, for tree=fill=cyan!25 [Chargrid(Katti et al., 2018); CUTIE(Zhao et al., 2019);BERTgrid(Denk and Reisswig, 2019); ACP(Palm et al., 2019);XYLayoutLM(Gu et al., 2022), for tree=fill=cyan!10] ] [Joint-learning Frameworks, for tree=fill=cyan!25 [TRIE(Zhang et al., 2020); VIES(Wang et al., 2021a), for tree=fill=cyan!10] ] [Relation-aware Models, for tree=fill=cyan!25 [Majumder et al. (Majumder et al., 2020a); Liu et al. (Liu et al., 2019a); PICK(Yu et al., 2021); FormNet(Lee et al., 2022b); FormNetv2(Lee et al., 2023), for tree=fill=cyan!10] ] [Few/Zero-shot Learning Models, for tree=fill=cyan!25 [LASER(Wang and Shang, 2022); Chen et al.(Chen et al., 2023); Cheng et al.(Cheng et al., 2020); QueryForm(Wang et al., 2023a), for tree=fill=cyan!10] ] [Prompt-based Frameworks, for tree=fill=cyan!25 [GenKIE(Cao et al., 2023b); ICL-D3IE(He et al., 2023); LMDX(Perot et al., 2023), for tree=fill=cyan!10] ] ] [Entity Linking, for tree=fill=blue!40 [Entity-level Linking, for tree=fill=blue!20 [DocStruct(Wang et al., 2020b); Zhang et al.(Zhang et al., 2021); KVPFormer.(Hu et al., 2023), for tree=fill=blue!10] ] [Token-level Linking, for tree=fill=blue!20 [Carbonell et al. (Carbonell et al., 2021); SPADE (Hwang et al., 2021);DocTR(Liao et al., 2023), for tree=fill=blue!10] ] ] [Visual Question Answering, for tree=fill=teal!40 [Single Page Frameworks, for tree=fill=teal!25 [Please refer Section 3. , for tree=fill=teal!10] ] [Multi-Page Frameworks, for tree=fill=teal!25 [Hi-VT5(Tito et al., 2023); GRAM(Blau et al., 2024); Kang et al. (Kang et al., 2024), for tree=fill=teal!10] ] ] ]
@@
Figure 2. Mono-task visually rich document understanding models
3.1. Key Information Extraction
@@
Key Information Extraction (KIE), a typical natural language processing task, refers to the task of identifying and extracting crucial pieces of information from textual data. Unlike typical name entity recognition methods for addressing plain text, VRDs contain visually rich entities like tables and charts, as well as spatial and logical layout arrangements to enhance the challenge of extracting crucial information. Although plain-text pretrained language models, such as BERT (Devlin, 2018), RoBERTa (Liu et al., 2019b), and ALBERT (Lan et al., 2019), are widely used as solid baselines on many benchmark datasets, more recent works introduce layout-aware pretrained models such as LayoutLM families (Xu et al., 2020a, b; Huang et al., 2022), LiLT (Wang et al., 2022b), Bros (Hong et al., 2021) to enhance the document representation by leveraging visual and layout information and achieving SoTA performance on several downstream tasks (see Section 3). This section will mainly focus on models specifically proposed for the document KIE models or only evaluated on KIE benchmark datasets including FUNSD (Jaume et al., 2019), CORD (Park et al., 2019), SROIE (Huang et al., 2019), XFUND (Xu et al., 2021), etc. Based on innovative aspects, we categorise KIE frameworks into five types: Feature-driven models use multimodal cues for rich feature representation. Joint Learning frameworks integrate auxiliary tasks to enhance document representation. Relation-aware models leverage spatial or logical relations via graphs or masked attention mechanisms. Few/Zero-shot learning frameworks explore methods for extracting key information with minimal labelled data, often using transfer learning. Prompt-based frameworks use structured prompts to guide specific information extraction from pretrained models or LLM/MLLMs. These categories represent diverse approaches to improving Key Information Extraction (KIE) by leveraging specific model designs and learning strategies tailored to document understanding challenges.
3.1.1. Feature-driven Models
@@
In the initial phase, certain Recurrent Neural Network (RNN)-based models (Lample et al., 2016) were introduced, primarily focusing on key information extraction tasks from plain text. However, these approaches overlook the importance of visual cues and layout information. Hence, several multimodal frameworks with feature-driven designs have been proposed to generate more representative document representations.
@@
Chargrid (Katti et al., 2018) first mentioned the significance of 2D structure for document KIE and designed a character-box-based chargrid to convert the textual and 2D layout structure into coloured visual cues feed into CNN. Instead of using fine-grained character information, CUTIE (Zhao et al., 2019) and BERTgrid (Denk and Reisswig, 2019) utilise various word embedding methods with bounding box coordinates to persevere the layout structure. ACP (Palm et al., 2019) leverages attention mechanism and multi-aspect features, including visual, semantic (both character and word) and spatial, into dilated CNN for capturing both short and long-term dependencies of each word piece.
@@
Joint learning KIE frameworks are proposed to leverage multi-level features to mitigate the information gap between various focused tasks. TRIE (Zhang et al., 2020) firstly offers an end-to-end framework for simultaneously conducting Optical Character Recognition (OCR) and KIE. OCR module will generate multi-aspect features, including positional, visual and textual aspects. Adaptively trainable weighting mechanisms are adopted to generate the fused embedding, followed by a Bi-LSTM-based Entity Extraction Module to conduct the final prediction. VIES (Wang et al., 2021a) is introduced to use the multi-level cues of vision, position and text to generate more comprehensive representations. Both token and segment (entity) level positional and visual features are acquired from text detection modules, and dual-level textual features are gathered by text recognition brunch, fused by a self-attention-based fusion module for sequence labelling.
3.1.2. Relation-aware Models
Instead of only leveraging multimodal information, leveraging the inherent spatial or logical relations between multi-level components may lead to more robust and comprehensive document representations. Majumder et al. (Majumder et al., 2020a) propose a model which starts by generating candidates for each field using type-specific detectors and key phrases. The neural model scores these candidates by learning dense representations that consider both textual content and spatial positioning, allowing for accurate information extraction across different document templates by focusing on the candidates’ relevance to the fields. Graph-based frameworks are increasingly favoured for modelling document elements’ spatial and logical relationships. This trend involves meticulously defining distinct graph structures and employing graph convolution techniques to incorporate these relationships into feature representations seamlessly.
Liu et al. (Liu et al., 2019a) first utilise a graph convolution module on top of the BiLSTM-CRF for key information extraction. Each document is a fully connected graph where the node is the textual representation $T$ of the document entity $E_{i}\in\mathbb{E}$, and the edge is related to the spatial relation between $E_{i}$ and another entity $E_{j}$. It is defined as $e_{ij}=[x_{ij},y_{ij},\frac{w_{i}}{h_{i}},\frac{h_{j}}{h_{i}},\frac{w_{j}}{h_{i}}]$, where $x_{ij}$ and $y_{ij}$ are horizontal and vertical distance between $E_{i}$ and $E_{j}$. The self-attention-based graph convolution is performed on node-edge-node triplets (concatenated nodes and edge embeddings) to learn contextually with all other nodes. PICK (Yu et al., 2021) adopts a similar way to construct the document graph but utilising multimodal node representations, including transformer-encoded textual embedding and CNN-encoded visual embedding, as well as the edge embedding is updated to $e_{ij}=[x_{ij},y_{ij},\frac{w_{i}}{h_{i}},\frac{h_{j}}{h_{i}},\frac{w_{j}}{h_{i}},\frac{S_{j}}{S_{i}}]$ where $S$ is the sentence length of corresponding entity. Additionally, to get the node embedding for the downstream tagging task, a soft adjacent matrix-based graph learning layer is applied to obtain the task-specific node representations.
FormNet (Lee et al., 2022b) introduces an end-to-end framework with a new attention mechanism, Rich Attention, and graph framework to make an order/distance sensitive long sequence transformer. In Rich Attention, the model introduces two order-based and pixel distance scores along the x/y axis. These are summed up with usual self-attention scores to enable the model order/distance awareness. Additionally, a Graph is applied to contextually learn the neighbouring token embedding before serialising the token sequence feeding into the transformer encoders. The graph nodes are text embedding of tokens, and the edges are relative positions between nodes. This could mitigate the imperfect serialization issue, enabling token representation capturing in more contexts. FormNetv2 (Lee et al., 2023) integrates image modality as the additional Graph edge feature to capture more visual cues  [^1] Please refer to Section 6.1.2 to get more details. To use contrastive loss to learn the multimodal graph representation, they first perform stochastic graph corruption to sample topology-corrupted and feature-corrupted graphs. Topology corruption randomly removes edges in the original graph, while feature corruption drops all modality features from nodes and edges. Then, the contrastive objective is to maximize agreement between a pair of tokens between corrupted and original graphs under a standard normalized temperature-scaled cross-entropy (Sohn, 2016) loss.
3.1.3. Few-shot Learning Frameworks
Extracting key information from VRDs based on deep learning frameworks typically involves more manual work to annotate the training data. However, acquiring large-scale, high-quality annotations in urgent and labouring-limited application scenarios is challenging. Thus, few-shot or one-shot frameworks emerged to extract the key information with less labouring annotations.
@@ 3.1.3. Few-shot Learning Frameworks
LASER (Wang and Shang, 2022) leverage the architecture from LayoutReader (Wang et al., 2021b) to reformulate the entity recognition task to sequence-labelling to generative model by embedding the entity type information (named label surface name) into the target sequence to enable the model label semantic-aware. LASER depends on a ”partially triangle” mask to use one encoder to encode source text and generate the prediction. The $n$ source text sequence, $\{t_{1},...,t_{n}\}$, is the text sequence by summing the relative word, spatial, and positional embedding, feeding into the $\mathcal{E}$ with full self-attention. The target sequence only attends previous tokens. The generative formulation is defined as:
(1) $t_{i-1},[B],t_{i},...,t_{j},[E],e_{1},...,e_{k},[T],t_{j+1}$
where $[B]$ and $[E]$ denote start and end of entity, $e_{i}$ are the label surface name; $[T]$ denotes the end of the label surface name. As label surface names and functional tokens do not exist, the learnable embeddings are applied to them to ensure the generative-progress layout-aware. A binary classification module is applied to classify the next generated token type of current generated token hidden states: from source or not, effectively controlling the generative series.
@@
Chen et al. (Chen et al., 2023) introduce a novel framework for entity-level, N-way soft-K-shot extracting key information from VRDs, focusing on the challenges of extracting rare or unseen entities from documents with few examples. It leverages a meta-learning approach (Snell et al., 2017; Oreshkin et al., 2018), utilizing a hierarchical decoder and contrastive learning (ContrastProtoNet) for task personalization and improved adaptation to new entity types. Additionally, it introduces FewVEX, a dataset tailored for entity-level few-shot VDER, to facilitate research and benchmarking in this area. The framework’s effectiveness is demonstrated through significant improvements in robustness and performance over existing meta-learning baselines.
In the domain of one-shot scenarios, Cheng et al. (Cheng et al., 2020) presents a novel application of graphs. They utilize attention mechanisms to seamlessly transfer spatial relationships between static text regions ( key/landmark) and dynamic text regions (value/field) from support documents to query documents. This transfer enables the acquisition of label probability distributions for files. Additionally, a self-attention-based module is integrated to exploit relationships between field entities, facilitating the derivation of label transition probability distributions. Finally, belief propagation is harnessed for inference within a pairwise Conditional Random Field (CRF), resulting in an assessable end-to-end trainable pipeline.
3.1.4. Prompt-learning Frameworks
QueryForm (Wang et al., 2023a) introduce a query-based framework for zero-shot document key information extraction. A dual prompting-based mechanism, entity query (E-prompt) and schema query (S-prompt), is introduced for pre-training and transferring knowledge from large-scale weakly annotated pre-trained webpages to target domains. The pretraining target is highly aligned with fine-tuning to ensure the proposed framework can consistently make query-conditional predictions at both stages. The HTML tags acquire the E-prompt ($\bm{e_{p}}$) during pretraining, and the S-prompt ($\bm{\tilde{s}_{p}}$) is generated from webpage domains, while during the fine-tuning, $\bm{e_{p}}$ is predefined and S-prompt $\bm{s_{p}}$ is learnable vectors. Supposing $\bm{t}$ is the serialised text of the input document, the pre-training and fine-tuning target $\bm{\hat{y}}$ and $\bm{y}$ can be represented as:
(2) $\bm{\hat{y}}=\mathcal{F}([\bm{s_{p}};\mathcal{E}[\bm{e_{p}},\bm{x}]]),$
(3) $\bm{y}=\mathcal{F}([\mathcal{E}[\bm{\tilde{s}_{p}};\bm{e_{p}},\bm{x}]])$
where $\mathcal{E}$ and $\mathcal{F}$ in here are the feature encoder and the rest of the language model, respectively. The training objective is to minimise the cross-entropy loss between $\bm{\hat{y}}$ and $\bm{y}$.
@@ 3.1.4. Prompt-learning Frameworks
Prompt learning is a technique in natural language processing where models are guided by specific prompts to produce or interpret targeted responses. With the rise of large-scale models, prompt learning can leverage contextual representation and implicit knowledge to enhance performance on targeted tasks. Since most LLMs (OpenAI, 2023; Touvron et al., 2023) and MLLMs (Liu et al., 2024a) are trained on plain text or natural images, layout-aware prompting and in-context learning (Brown et al., 2020) methods have been introduced to improve understanding in visually rich documents (VRD)
@@
GenKIE (Cao et al., 2023b) proposes an encoder-decoder-based multimodal KIE framework to leverage prompt to adapt various datasets and better leverage multimodal information. Following (Xu et al., 2020a, b), different encoding methods are adopted to acquire textual, layout and visual embeddings. Byte Pair Encoding is used as language pretrained backbones, and the OCR extracted document content is concatenated with predefined prompts split by the ”[SEP]” token between OCR tokens and each prompt. The 2D-positional encoding introduced by (Xu et al., 2020a) is used to acquire the layout embedding of each OCR extracted token, and ResNet (He et al., 2016) is used to extract the visual representations following (Wang et al., 2022c). The multimodal representations are fed into an encoder to learn interactively between modalities. Prompts, inserted at the end of the encoder’s textual inputs, are either template-style or question-style. For the entity extraction task, the prompt specifies the target entity type, and the decoder generates the entity value (e.g., for ”Company is?”, the decoder outputs the company name). For the entity-labeling task, the prompt includes the value, and the decoder provides the entity type (e.g., for ”Es Kopi Rupa is [SEP]”, the decoder identifies the entity type).
@@
ICL-D3IE (He et al., 2023) is the first framework to employ LLMs with in-context learning to extract key information from VRDs using the iterative updated diverse demonstrates. Before designing the initial diverse demonstrations, the most similar $n$ training documents to the $n$ test samples need to be selected by calculating the cosine-similarity of document representations encoded by Sentence-BERT (Reimers and Gurevych, 2019). Then, different types of demonstrations are introduced to integrate multiple-view context information into LLMs. Hard Demonstrations highlight the most challenging cases, are initially designed based on the incorrectly predicted cases from GPT-3 (Brown et al., 2020) predictions and are updated based on prediction results during the training process. Layout-aware demonstrations are created by selecting the adjacent hard segments to understand the positional relations. Formatting demonstrations are designed to guide LLMs in formatting the outputs for easy post-processing.
@@
LMDX (Perot et al., 2023) designs a pipeline to use arbitrary LLMs to extract singular, repeated and hierarchical entities from VRDs. The document images are first fed into off-the-shelf OCR tools and are divided into smaller document chunks to be processed by the LLMs with accessible input length. Then, prompts are generated under XML-like tags to control the LLM’s responses and mitigate hallucination. Document Representation is a prompt contains the chunk content with the coordinates of OCR lines to bring layout modality to LLMs, where a text segment extracted by OCR tools is represented ¡text¿ [$x_{centre}$, $y_{centre}$]. After that, the task description and scheme representation prompts are designed to explain the task to accomplish and determine the output format. During inference, $N$ prompts with $K$ LLMs completions are generated to sample the correct answers.
3.1.5. Summary of Key Information Extraction

### v2

@@ 3 Mono-Task Document Understanding Frameworks
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.8cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2 [Mono-task based models (Content Understanding), for tree=fill=lime!60 [Key Information Extraction, for tree=fill=cyan!40 [Feature-driven Models, for tree=fill=cyan!25 [Chargrid[chargrid]; CUTIE[cutie]; BERTgrid[bertgrid]; ACP[palm2019attend]; XYLayoutLM[xylayoutlm], for tree=fill=cyan!10] ] [Joint-learning Frameworks, for tree=fill=cyan!25 [TRIE[trie]; VIES[ephoie], for tree=fill=cyan!10] ] [Relation-aware Models, for tree=fill=cyan!25 [Majumder et al. [majumder2020representation]; Liu et al. [liu2019graph]; PICK[pick]; FormNet[formnet]; FormNetv2[formnetv2], for tree=fill=cyan!10] ] [Few/Zero-shot Learning Models, for tree=fill=cyan!25 [LASER[laser]; Chen et al.[chen2023task]; Cheng et al.[cheng2020one]; QueryForm[queryform], for tree=fill=cyan!10] ] [Prompt-based Frameworks, for tree=fill=cyan!25 [GenKIE[genkie]; ICL-D3IE[icld3ie]; LMDX[lmdx], for tree=fill=cyan!10] ] ] [Entity Linking, for tree=fill=blue!40 [Entity-level Linking, for tree=fill=blue!20 [DocStruct[docstruct]; Zhang et al.[zhang2021entity]; KVPFormer.[kvpformer], for tree=fill=blue!10] ] [Word-level Linking, for tree=fill=blue!20 [Carbonell et al. [carbonell2021named]; SPADE [hwang2021spatial];DocTR[doctr], for tree=fill=blue!10] ] ] [Question Answering, for tree=fill=teal!40 [Single Page Frameworks, for tree=fill=teal!25 [Please refer to Section 3. , for tree=fill=teal!10] ] [Multi-Page Frameworks, for tree=fill=teal!25 [Hi-VT5[mpdocvqa]; GRAM[gram]; Kang et al. [kang2024multi], for tree=fill=teal!10] ] ] ]
@@
Figure 2: Mono-task visually rich document understanding models
3.1 Key Information Extraction
@@
Key Information Extraction (KIE), a typical natural language processing task, refers to the task of identifying and extracting crucial pieces of information from textual data. Unlike typical name entity recognition methods for addressing plain text, VRDs contain visually rich entities like tables and charts, as well as spatial and logical layout arrangements to enhance the challenge of extracting crucial information. Although plain-text pretrained language models, such as BERT [bert], RoBERTa [roberta], and ALBERT [albert], are widely used as solid baselines on many benchmark datasets, more recent works introduce layout-aware pretrained models such as LayoutLM families [layoutlm, layoutlmv2, layoutlmv3], LiLT [lilt], Bros [bros] to enhance the document representation by leveraging visual and layout information and achieving SoTA performance on several downstream tasks (see Section 3). This section will mainly focus on models specifically proposed for the document KIE models or only evaluated on KIE benchmark datasets including FUNSD [funsd], CORD [cord], SROIE [sroie], XFUND [layoutxlm], etc. Based on innovative aspects, we categorise KIE frameworks into five types: Feature-driven models use multimodal cues for rich feature representation. Joint Learning frameworks integrate auxiliary tasks to enhance document representation. Relation-aware models leverage spatial or logical relations via graphs or masked attention mechanisms. Few/Zero-shot learning frameworks explore methods for extracting key information with minimal labelled data, often using transfer learning. Prompt-based frameworks use structured prompts to guide specific information extraction from pretrained models or LLM/MLLMs. These categories represent diverse approaches to improving KIE by leveraging specific model designs and learning strategies tailored to document understanding challenges.
3.1.1 Feature-driven Models
@@
In the initial phase, certain Recurrent Neural Network (RNN)-based models [lample2016neural] were introduced, primarily focusing on key information extraction tasks from plain text. However, these approaches overlook the importance of visual cues and layout information. Hence, several multimodal frameworks with feature-driven designs have been proposed to generate more representative document representations.
@@
Chargrid [chargrid] first mentioned the significance of 2D structure for document KIE and designed a character-box-based chargrid to convert the textual and 2D layout structure into coloured visual cues feed into CNN. Instead of using fine-grained character information, CUTIE [cutie] and BERTgrid [bertgrid] utilise various word embedding methods with bounding box coordinates to persevere the layout structure. ACP [palm2019attend] leverages attention mechanism and multi-aspect features, including visual, semantic (both character and word) and spatial, into dilated CNN for capturing both short and long-term dependencies of each word piece.
@@
Joint learning KIE frameworks are proposed to leverage multi-level features to mitigate the information gap between various focused tasks. TRIE [trie] firstly offers an end-to-end framework for simultaneously conducting Optical Character Recognition (OCR) and KIE. OCR module will generate multi-aspect features, including positional, visual and textual aspects. Adaptively trainable weighting mechanisms are adopted to generate the fused embedding, followed by a Bi-LSTM-based Entity Extraction Module to conduct the final prediction. VIES [ephoie] is introduced to use the multi-level cues of vision, position and text to generate more comprehensive representations. Both token and segment (entity) level positional and visual features are acquired from text detection modules, and dual-level textual features are gathered by text recognition brunch, fused by a self-attention-based fusion module for sequence labelling.
3.1.2 Relation-aware Models
Beyond leveraging multimodal features, modeling the spatial and logical relationships between document components can yield more robust and context-aware representations. Relation-aware models explicitly encode these inter-element dependencies—often using graph structures—to support complex information extraction tasks. Majumder et al. [majumder2020representation] propose a two-stage approach: field-specific detectors first generate candidates, which are then scored by a neural model that encodes both textual and spatial features. This enables the model to adapt across varying document layouts by focusing on field relevance.
3.1.3 Relation-aware Models
Beyond leveraging multimodal signals, modeling the spatial and logical relationships between document components can yield more robust and context-aware representations. Relation-aware models explicitly encode these inter-element dependencies—often using graph structures—to support complex information extraction tasks. Majumder et al. [majumder2020representation] propose a two-stage approach: field-specific detectors first generate candidates, which are then scored by a neural model that encodes both textual and spatial features. This enables the model to adapt across varying document layouts by focusing on field relevance.
Graph-based frameworks have gained traction for encoding spatial/logical relationships. Liu et al. [liu2019graph] pioneer this by constructing a fully connected graph where each node is a textual entity $E_{i}\in\mathbb{E}$ and edges capture relative spatial features:
$e_{ij}=\left[x_{ij},\ y_{ij},\ \frac{w_{i}}{h_{i}},\ \frac{h_{j}}{h_{i}},\ \frac{w_{j}}{h_{i}}\right]$
Graph convolution is then applied over node-edge-node triplets using self-attention to propagate contextual information.
PICK [pick] extends this design with multimodal node features, combining transformer-based text embeddings and CNN-based visual features. The edge representation is also enriched with sentence length information. To enhance tagging performance, PICK applies a soft adjacency matrix for graph learning, yielding task-specific node embeddings.
FormNet [formnet] introduces a novel attention mechanism—Rich Attention—that incorporates positional order and pixel distances (along $x/y$ axes) into self-attention scores. This allows transformers to be aware of layout order and spatial proximity. Additionally, a graph encoder integrates neighborhood context prior to serialization, mitigating order ambiguity.
FormNetv2 [formnetv2] further enhances this by introducing visual cues as edge features and applying contrastive learning. Two types of graph corruption—topological (edge dropping) and feature-level (modality masking)—are used to generate augmented views. The model then learns robust multimodal embeddings using a normalized temperature-scaled cross-entropy loss [sohn2016improved].
In summary, these relation-aware models highlight the shift from flat sequence modeling to structure-aware document representation. While their specific graph construction and fusion techniques vary, they all aim to capture the rich spatial and semantic dependencies crucial for key information extraction and related tasks.
3.1.4 Few-shot Learning Frameworks
Extracting key information from VRDs using deep learning typically requires extensive manual annotation. However, in many real-world scenarios, obtaining large-scale, high-quality labels is costly and time-consuming. To address this challenge, few-shot and one-shot learning frameworks have been developed to enable key information extraction with minimal annotated data.
@@ 3.1.4 Few-shot Learning Frameworks
LASER [laser] builds upon the architecture of LayoutReader [layoutreader], reformulating entity recognition from a sequence labeling task into a generative one. By embedding entity type information—referred to as label surface names—into the target sequence, LASER enables the model to become label-semantic aware. It uses a Partially Triangular attention mask, allowing a single encoder to encode both the source text and the generated sequence. The input sequence $\{t_{1},...,t_{n}\}$, comprising the sum of word, spatial, and positional embeddings, is fed into the encoder $\mathcal{E}$ with full self-attention. The decoder attends only to previously generated tokens. The generative target format is defined as:
$t_{i-1},\ [B],\ t_{i},...,t_{j},\ [E],\ e_{1},...,e_{k},\ [T],\ t_{j+1}$ (1)
where $[B]$ and $[E]$ denote the entity span, $e_{i}$ represent the label surface name, and $[T]$ marks the end of the label name. Since label tokens and special symbols do not appear in the source, learnable embeddings are assigned to them. A binary classifier predicts whether the next token should be copied from the source, effectively guiding the generative process.
@@
Chen et al. [chen2023task] propose an entity-level N-way soft-K-shot learning framework for key information extraction under few-shot settings. Targeting rare or unseen entity types, their approach employs meta-learning [snell2017prototypical, oreshkin2018tadam] with a hierarchical decoder and contrastive learning module (ContrastProtoNet) for task personalization and improved adaptation. They also introduce FewVEX, a benchmark dataset designed for entity-level few-shot VRD extraction, and demonstrate that their method significantly outperforms existing meta-learning baselines in robustness and performance.
In the context of one-shot learning, Cheng et al. [cheng2020one] propose a graph-based method that transfers spatial relationships from support to query documents. Their approach leverages attention mechanisms to model interactions between static regions (keys/landmarks) and dynamic regions (values/fields), enabling accurate label probability distribution inference. A self-attention module captures dependencies among field entities, while a pairwise Conditional Random Field (CRF) with belief propagation completes the inference, resulting in an end-to-end trainable pipeline.
3.1.5 Prompt-learning Frameworks
@@
Prompt learning is a technique in natural language processing that guides models using specific prompts to elicit targeted responses. With the emergence of large-scale models, it has become an effective way to leverage contextual representations and the implicit knowledge embedded within pretrained models. Since most LLMs [chatgpt, llama] and MLLMs [llava] are primarily trained on plain text or natural images, layout-aware prompting and in-context learning methods [gpt3] have been proposed to adapt these models for VRD-CU.
QueryForm [queryform] introduces a query-based framework for zero-shot document key information extraction. It employs a dual prompting mechanism—entity prompts (E-prompts) and schema prompts (S-prompts)—to transfer knowledge from large-scale, weakly annotated pretraining data (e.g., webpages) to target domains. The design ensures that the pretraining and fine-tuning objectives are well aligned, enabling consistent query-conditional predictions across both stages. During pretraining, HTML tags are used to construct the E-prompt $\bm{e_{p}}$, while the S-prompt $\bm{\tilde{s}_{p}}$ is derived from webpage domain information. In the fine-tuning stage, $\bm{e_{p}}$ becomes predefined, and $\bm{s_{p}}$ is treated as a set of learnable vectors. Let $\bm{x}$ denote the serialized text of the input document. The prediction targets during pretraining and fine-tuning are defined as:
$\bm{\hat{y}}$ $=\mathcal{F}([\bm{s_{p}};\ \mathcal{E}[\bm{e_{p}},\bm{x}]]),$ (2)
$\bm{y}$ $=\mathcal{F}([\mathcal{E}[\bm{\tilde{s}_{p}};\ \bm{e_{p}},\bm{x}]])$ (3)
where $\mathcal{E}$ is the feature encoder and $\mathcal{F}$ represents the remaining layers of the language model. The training objective is to minimize the cross-entropy loss between $\bm{\hat{y}}$ and $\bm{y}$.
@@
GenKIE [genkie] proposes an encoder-decoder-based multimodal KIE framework to leverage prompt to adapt various datasets and better leverage multimodal information. Following [layoutlm, layoutlmv2], different encoding methods are adopted to acquire textual, layout and visual embeddings. Byte Pair Encoding is used as language pretrained backbones, and the OCR extracted document content is concatenated with predefined prompts split by the ”[SEP]” token between OCR tokens and each prompt. The 2D-positional encoding introduced by [layoutlm] is used to acquire the layout embedding of each OCR extracted token, and ResNet [resnet] is used to extract the visual representations following [wang2022ofa]. The multimodal representations are fed into an encoder to learn interactively between modalities. Prompts, inserted at the end of the encoder’s textual inputs, are either template-style or question-style. For the entity extraction task, the prompt specifies the target entity type, and the decoder generates the entity value (e.g., for ”Company is?”, the decoder outputs the company name). For the entity-labeling task, the prompt includes the value, and the decoder provides the entity type (e.g., for ”Es Kopi Rupa is [SEP]”, the decoder identifies the entity type).
@@
ICL-D3IE [icld3ie] is the first framework to employ LLMs with in-context learning to extract key information from VRDs using the iterative updated diverse demonstrates. Before designing the initial diverse demonstrations, the most similar $n$ training documents to the $n$ test samples need to be selected by calculating the cosine-similarity of document representations encoded by Sentence-BERT [sentencebert]. Then, different types of demonstrations are introduced to integrate multiple-view context information into LLMs. Hard Demonstrations highlight the most challenging cases, are initially designed based on the incorrectly predicted cases from GPT-3 [gpt3] predictions and are updated based on prediction results during the training process. Layout-aware demonstrations are created by selecting the adjacent hard segments to understand the positional relations. Formatting demonstrations are designed to guide LLMs in formatting the outputs for easy post-processing.
@@
LMDX [lmdx] designs a pipeline to use arbitrary LLMs to extract singular, repeated and hierarchical entities from VRDs. The document images are first fed into off-the-shelf OCR tools and are divided into smaller document chunks to be processed by the LLMs with accessible input length. Then, prompts are generated under XML-like tags to control the LLM’s responses and mitigate hallucination. Document Representation is a prompt contains the chunk content with the coordinates of OCR lines to bring layout modality to LLMs. After that, the task description and scheme representation prompts are designed to explain the task to accomplish and determine the output format. During inference, $N$ prompts with $K$ LLMs completions are generated to sample the correct answers.
3.1.6 Summary of Key Information Extraction